A body is falling free in a vacuum. The fall is necessarily related to the gravitational

acceleration g and the height h from which the body is dropped. Use dimensional

analysis to show that the velocity V of the falling body satisfies the relation V/√(gh)=constant.


1
Expert's answer
2021-06-01T08:50:31-0400

Dimension of V is "\dfrac{m}{sec}."


Dimension of g is "\dfrac{m}{sec^2}."


Dimension of h is "m."


Thus, dimension of "\sqrt{gh}" is 

"\sqrt{m\cdot \frac{m}{sec^2}}=\dfrac{m}{sec}=\text{dimension of V.}"

 

Therefore, the relation "\dfrac{V}{\sqrt{gh}}=constant."